### v2

@@ 3.3 Affine boundary conditions
The objective of this section is to formulate $sl(2,\mathbb{R})$ higher–spin $AdS_{2}$ dilaton gravity under affine boundary conditions. Our analysis proceeds by computing the asymptotic symmetry algebra associated with the most general (i.e., least restrictive) set of boundary data. We begin by introducing the $a_{t}$ component of the gauge connection, which takes values in the $sl(2,\mathbb{R})$ Lie algebra:
@@
which defines an additional set of three state–dependent functions ${\cal X}^{i}$. Our aim is to derive the resulting asymptotic symmetry algebra under affine boundary conditions through a canonical analysis. For this purpose, we systematically investigate all gauge transformations of the form (14) that preserve the boundary data encoded in equations (20) and (49).
At this stage, it is convenient to parameterize the gauge transformation in terms of the $sl(2,\mathbb{R})$ Lie algebra basis:
@@
The infinitesimal variation of these charges, which encodes the structure of the asymptotic symmetry algebra, is given by Banados:1994tn :
@@
Having established both the infinitesimal gauge transformations and the associated canonical boundary charges, we now derive the asymptotic symmetry algebra using the standard procedure outlined in Blagojevic:2002du . The algebraic structure is encoded in the following relation:
@@
where $\tau_{12}=\tau_{1}-\tau_{2}$. The coefficients $\eta^{ij}=\mathfrak{tr}({\tt L}_{i}{\tt L}_{j})$ define the components of the invariant bilinear form in the fundamental representation of the $sl(2,\mathbb{R})$ Lie algebra. An alternative and more compact representation of the operator product algebra can be written using index-free notation as follows:
@@
Here, $\eta^{AB}$ denotes the trace matrix, and $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ are the structure constants of the underlying algebra, with indices running over $(A,B=0,\pm 1)$. In particular, one has $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Accordingly, under the most relaxed set of boundary conditions for $sl(2,\mathbb{R})$ dilaton gravity, the resulting asymptotic symmetry algebra is identified with a single copy of the affine $\mathfrak{sl}(2,\mathbb{R})_{k}$ algebra.
The time–dependent profile of the dilaton field at the affine boundary introduces structural modifications that go beyond the classical asymptotic symmetry algebra of $sl(2,\mathbb{R})$ JT dilaton gravity. In the standard BF formulation, the asymptotic symmetry algebra is determined solely by gauge transformations of the $sl(2,\mathbb{R})$–valued connection $a_{t}$. However, once the $sl(2,\mathbb{R})$–valued dilaton field $x$ is treated as a dynamical degree of freedom, the set of residual symmetries becomes further constrained, giving rise to an enriched physical structure. In this extended framework, the symmetry algebra must incorporate not only the reparametrization modes but also new dynamical currents induced by the dilaton.
As a consequence, the full symmetry algebra necessitates the inclusion of abelian currents alongside the classical $\mathfrak{sl}(2,\mathbb{R})_{k}$ affine modes. This reveals that the affine boundary dynamics of JT gravity cannot be adequately described by the Schwarzian action alone, but instead call for a more general, multi-mode effective action. In this refined picture, the dilaton field plays a dual role: it determines both the geometric background and the extended algebraic structure realized at the boundary.
@@ 3.4 Conformal boundary conditions
In this section, our goal is to analyze the asymptotic symmetry algebra associated with Brown–Henneaux-type boundary conditions. This gauge choice is specifically designed to make the asymptotic symmetry structure manifest and to streamline the analysis of boundary dynamics. By decoupling the radial dependence, it effectively isolates the relevant asymptotic degrees of freedom, thereby facilitating the computation of conserved quantities and the identification of corresponding boundary theories.
This gauge plays a central role in maintaining the full asymptotic symmetry group and in determining the effective boundary action, particularly within the context of holographic duality. To proceed, we impose the Drinfeld–Sokolov highest weight gauge condition on the $sl(2,\mathbb{R})$-valued connection defined in Eq. (20), which further restricts the allowed functional form of the coefficients. Under this reduction, the fields are constrained as follows:
@@
Here, $\alpha$ acts as a scaling parameter whose exact value will be fixed at a later stage. We define two key functionals: the boundary charge $\mathcal{L}$ and the dilaton field $\cal X$}. With these constraints in place, we are now prepared to derive the conformal asymptotic symmetry algebra. As dictated by the Drinfeld–Sokolov reduction, the residual gauge parameter $\lambda$ depends on four independent functions, with $\epsilon\equiv\epsilon^{+1}$ serving as the primary functional degree of freedom. The gauge parameter then takes the following form:
@@
Under the imposed boundary conditions, the corresponding canonical charges $\mathcal{Q}_{a}[\lambda]$ and $\mathcal{Q}_{x}[\lambda]$ remain integrable, thereby allowing $\mathcal{L}$ and $\mathcal{X}$ to be interpreted as conserved quantities. These gauge transformations are instrumental in uncovering the structure of the asymptotic symmetry algebra Banados:1998pi .
@@
From these relations, it follows that $\cal X$} transforms as a vector under boundary diffeomorphisms, while $\mathcal{L}$ behaves as a spin-two density. Throughout this analysis, the explicit dependence on the gauge parameter $\lambda$ has been omitted, as the boundary conditions no longer require it. In effect, the residual symmetries associated with the connection $a_{t}$ and the dilaton $x$ are interpreted as diffeomorphisms on the boundary circle, generated by the vector field $\epsilon$. It is thus observed that the asymptotic symmetry algebra of $sl(2,\mathbb{R})$ dilaton gravity with conformal boundary conditions is given by a single copy of the Virasoro algebra with central charge $c=3k$.
The conformal boundary profile of the dilaton field induces a structural extension analogous to the affine case, leading to an enriched version of the classical asymptotic symmetry algebra in $sl(2,\mathbb{R})$ JT gravity. In the conventional BF formulation, the asymptotic symmetry algebra is generated solely by gauge transformations of the $sl(2,\mathbb{R})$-valued connection $a_{t}$. However, once the dilaton field $x$ is promoted to a dynamical quantity, the residual symmetry subalgebra becomes more constrained, resulting in a more intricate physical structure. This generalized framework requires that the full symmetry algebra incorporate both the standard reparametrization modes and additional currents sourced directly by the dynamical behavior of the dilaton.
As a consequence, the boundary symmetry structure must necessarily include abelian extensions beyond the classical Virasoro algebra. This observation suggests that the Schwarzian action alone is insufficient to capture the complete boundary dynamics in the presence of a nontrivial dilaton configuration. Instead, a multi–component extended action is required to account for all dynamical features. In this context, the dilaton field assumes a dual role, encoding not only the background geometry but also the full algebraic structure realized at the boundary.
@@ 4.1 $\mathcal{N}=1$ affine boundary conditions
The goal of this section is to construct an extended higher–spin $AdS_{2}$ framework by formulating it as an $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ dilaton supergravity theory defined with affine boundary conditions. Our analysis aims to determine the resulting asymptotic symmetry algebra under the most general boundary conditions consistent with this setup.
Following the discussion in the previous section, we adopt the principal embedding of $sl(2,\mathbb{R})$ into $\mathfrak{osp}(1|2)$ as a subalgebra. This embedding provides a natural foundation for implementing affine boundary conditions in asymptotically $AdS_{2}$ spacetimes. To proceed with this construction, we introduce the gauge connection in the following form:
@@
In this framework, we obtain eight additional state-dependent functions, ${\cal X}^{i}$ and ${\cal Y}^{i}$, arising from the dilaton field expansion. Our goal is to derive the asymptotic symmetry algebra associated with these affine boundary conditions via a canonical analysis. To this end, we systematically investigate all gauge transformations that preserve the structure of the boundary conditions defined above.
@@
As a final step in this construction, we define the canonical boundary charge $\mathcal{Q}_{a,x}[\lambda]$, which generates the gauge transformations presented in Eqs. (51)–(56). The variation of this charge, which encodes the structure of the asymptotic symmetry algebra, is given by Banados:1994tn :
@@
Having established both the infinitesimal gauge transformations and the canonical boundary charge, we are now in a position to derive the asymptotic symmetry algebra using the standard procedure outlined in Blagojevic:2002du . This is accomplished through the fundamental relation:
@@
Here, $\eta^{AB}$ denotes the invariant trace matrix, while $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ represent the structure constants of the $\mathfrak{osp}(1|2)$ algebra, with indices $(A,B=0,\pm 1,\pm\tfrac{1}{2})$. These constants satisfy the relation $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Under the most general (least restrictive) boundary conditions, the asymptotic symmetry algebra of $\mathfrak{osp}(1|2)$ dilaton supergravity is characterized by a single copy of the affine $\mathfrak{osp}(1|2)_{k}$ algebra.
In the context of $\mathfrak{osp}(1|2)$ JT dilaton gravity with affine boundary conditions, much like in the $sl(2,\mathbb{R})$ JT dilaton gravity under analogous assumptions, the asymptotic symmetry algebra is initially given by the infinite–dimensional $\mathfrak{osp}(1|2)_{k}$ algebra. However, the presence of a time-dependent dilaton field dynamically deforms this structure. As a result, even the expected preservation of the finite $\mathfrak{osp}(1|2)$ subalgebra within $\mathfrak{osp}(1|2)_{k}$ is partially broken. The dilaton acts as a stabilizing agent for the gauge superconnection and, through its boundary fluctuations, obstructs the full realization of the infinite–dimensional affine symmetry.
This symmetry breaking is attributed to the influence of the dilaton on free boundary excitations. Consequently, the residual symmetries encoded in the fields $a_{t}$ and $x$ correspond to boundary diffeomorphisms, generated respectively by the bosonic and fermionic components $\epsilon^{i}$ and $\eta^{i}$ of the gauge parameter.
@@ 4.2 $\mathcal{N}=1$ superconformal boundary conditions
In this section, we aim to analyze the asymptotic symmetry algebra associated with the Brown–Henneaux boundary conditions in the context of $\mathfrak{osp}(1|2)$ supergravity. To this end, we impose the Drinfeld–Sokolov highest weight gauge condition on the $\mathfrak{osp}(1|2)$-valued connection defined in Eq. (48), which further restricts the coefficient structure. This gauge fixing yields the following field constraints:
@@
After implementing the above procedures, we are now positioned to derive the conformal asymptotic symmetry algebra. As a consequence of the Drinfeld–Sokolov reduction, the residual gauge parameter $\lambda$ is effectively described by only four independent functions. These are denoted by $\epsilon\equiv\epsilon^{+1}$ and $\eta\equiv\eta^{+\frac{1}{2}}$, and are explicitly given by:
@@
Importantly, these gauge transformations provide key insights into the structure of the asymptotic symmetry algebra Banados:1998pi . By analyzing the behavior of these asymptotic symmetries, one can determine the integral form of the variation of the canonical boundary charges,
@@
Table 1: superconformal spins for superconformal charges and related dilatons.
In the context of $sl(2,\mathbb{R})$ JT dilaton gravity, as well as in its $\mathfrak{osp}(1|2)$ extension under conformal boundary conditions, the asymptotic symmetry algebra at the boundary is initially identified with the infinite–dimensional $\mathcal{N}=1$ superconformal algebra. However, the presence of a time–dependent dilaton field alters the algebraic structure of the asymptotic symmetry. As a result, even the expected preservation of the finite $\tt{O}\tt{S}p(1|2)$ subgroup within the full $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ symmetry is dynamically reduced, and the remaining infinite–dimensional symmetry components are broken.
@@
This choice retains only the $\mathcal{X}^{0}$ component while suppressing all fermionic modes $\mathcal{Y}^{\pm\frac{1}{2}}$ and the bosonic charges $\mathcal{X}^{\pm 1}$. Under the gauge transformations generated by the full affine algebra (Eqs. (51)–(56)), one finds that such a configuration leads to a restricted set of residual symmetries. In particular, the variations $\delta_{\lambda}\mathcal{Y}^{\pm\frac{1}{2}}$ and $\delta_{\lambda}\mathcal{X}^{\pm 1}$ vanish identically, while $\delta_{\lambda}\mathcal{X}^{0}$ survives only for $\epsilon^{0}$ and a subset of $\eta$–parameters. As a consequence, the boundary charges associated with broken modes cease to transform non–trivially and are effectively projected out of the asymptotic symmetry algebra. This demonstrates, at a minimal level, how specific dilaton backgrounds can dynamically suppress certain symmetry generators, resulting in a reduced realization of the full $\mathfrak{osp}(1|2)_{k}$ structure. This kind of toy model construction, where a specific field configuration is used to analyze residual symmetry algebras, is also employed in the context of AdS2 boundary conditions Grumiller:2017qao .
@@
We therefore identify the asymptotic symmetry algebra as the classical extended $\mathcal{N}=1$ superconformal algebra, rather than its quantum counterpart. This classical algebra arises naturally in our model and is consistent with the symmetry–breaking mechanism induced by the dilaton. Up to this point, our discussion has remained within the classical regime of the extended $\mathcal{N}=1$ superconformal symmetry. The quantum version of the theory, however, has already been presented in Eqs. (4)–(10). It is natural to ask whether the quantum theory can also be formulated in terms of a classical extended superconformal symmetry, or whether it necessitates a distinct quantum realization. While this remains a nontrivial question, preliminary evidence suggests that the quantum structure may be realized consistently, preserving the essential features of the classical algebra.

### v3

@@ 3.3 Affine boundary conditions
The objective of this section is to formulate $\mathfrak{sl}(2,\mathbb{R})$ higher–spin $AdS_{2}$ dilaton gravity under affine boundary conditions. Our analysis proceeds by computing the ASA associated with the most general (i.e., least restrictive) set of boundary data. We begin by introducing the $a_{t}$ component of the gauge connection, which takes values in the $\mathfrak{sl}(2,\mathbb{R})$ Lie algebra:
@@
which defines an additional set of three state–dependent functions ${\cal X}^{i}$. Our aim is to derive the resulting ASA under affine boundary conditions through a canonical analysis. For this purpose, we systematically investigate all gauge transformations of the form (14) that preserve the boundary data encoded in equations (20) and (49).
At this stage, it is convenient to parameterize the gauge transformation in terms of the $\mathfrak{sl}(2,\mathbb{R})$ Lie algebra basis:
@@
The infinitesimal variation of these charges, which encodes the structure of the ASA, is given by Banados:1994tn :
@@
Having established both the infinitesimal gauge transformations and the associated canonical boundary charges, we now derive the ASA using the standard procedure outlined in Blagojevic:2002du . The algebraic structure is encoded in the following relation:
@@
where $\tau_{12}=\tau_{1}-\tau_{2}$. The coefficients $\eta^{ij}=\mathfrak{tr}({\tt L}_{i}{\tt L}_{j})$ define the components of the invariant bilinear form in the fundamental representation of the $\mathfrak{sl}(2,\mathbb{R})$ Lie algebra. An alternative and more compact representation of the operator product algebra can be written using index-free notation as follows:
@@
Here, $\eta^{AB}$ denotes the trace matrix, and $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ are the structure constants of the underlying algebra, with indices running over $(A,B=0,\pm 1)$. In particular, one has $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Accordingly, under the most relaxed set of boundary conditions for $\mathfrak{sl}(2,\mathbb{R})$ dilaton gravity, the resulting ASA is identified with a single copy of the affine $\mathfrak{sl}(2,\mathbb{R})_{k}$ algebra.
The time–dependent profile of the dilaton field at the affine boundary does not modify the affine $\mathfrak{sl}(2,\mathbb{R})_{k}$ algebra itself. In the standard BF formulation, the ASA is determined by gauge transformations of the $\mathfrak{sl}(2,\mathbb{R})$–valued connection $a_{t}$, and this algebraic structure remains intact.
However, once the $\mathfrak{sl}(2,\mathbb{R})$–valued dilaton field $x$ is treated as a dynamical boundary degree of freedom, the set of background–preserving residual gauge transformations becomes restricted. This does not deform the affine algebra, but it affects how the symmetry is realized on the boundary phase space.
Consequently, while the ASA remains a single copy of affine $\mathfrak{sl}(2,\mathbb{R})_{k}$, the boundary dynamics is governed by an enlarged phase space that includes additional dynamical modes associated with the dilaton. In this sense, the Schwarzian action alone does not exhaust the full set of boundary degrees of freedom in the affine regime. The dilaton thus plays a dual role: it determines the geometric background and restricts the admissible residual gauge transformations to those that preserve a given configuration, while the affine algebra itself remains unchanged.
@@ 3.4 Conformal boundary conditions
In this section, our goal is to analyze the ASA associated with Brown–Henneaux-type boundary conditions. This gauge choice is specifically designed to make the asymptotic symmetry structure manifest and to streamline the analysis of boundary dynamics. By decoupling the radial dependence, it effectively isolates the relevant asymptotic degrees of freedom, thereby facilitating the computation of conserved quantities and the identification of corresponding boundary theories.
This gauge plays a central role in maintaining the full asymptotic symmetry group and in determining the effective boundary action, particularly within the context of holographic duality. To proceed, we impose the Drinfeld–Sokolov highest weight gauge condition on the $\mathfrak{sl}(2,\mathbb{R})$-valued connection defined in Eq. (20), which further restricts the allowed functional form of the coefficients. Under this reduction, the fields are constrained as follows:
@@
Here, $\alpha$ acts as a scaling parameter whose exact value will be fixed at a later stage. We define two key functionals: the boundary charge $\mathcal{L}$ and the dilaton field $\cal X$}. With these constraints in place, we are now prepared to derive the conformal ASA. As dictated by the Drinfeld–Sokolov reduction, the residual gauge parameter $\lambda$ depends on four independent functions, with $\epsilon\equiv\epsilon^{+1}$ serving as the primary functional degree of freedom. The gauge parameter then takes the following form:
@@
Under the imposed boundary conditions, the corresponding canonical charges $\mathcal{Q}_{a}[\lambda]$ and $\mathcal{Q}_{x}[\lambda]$ remain integrable, thereby allowing $\mathcal{L}$ and $\mathcal{X}$ to be interpreted as conserved quantities. These gauge transformations are instrumental in uncovering the structure of the ASA Banados:1998pi .
@@
From these relations, it follows that $\cal X$} transforms as a vector under boundary diffeomorphisms, while $\mathcal{L}$ behaves as a spin-two density. Throughout this analysis, the explicit dependence on the gauge parameter $\lambda$ has been omitted, as the boundary conditions no longer require it. In effect, the residual symmetries associated with the connection $a_{t}$ and the dilaton $x$ are interpreted as diffeomorphisms on the boundary circle, generated by the vector field $\epsilon$. It is thus observed that the ASA of $\mathfrak{sl}(2,\mathbb{R})$ dilaton gravity with conformal boundary conditions is given by a single copy of the Virasoro algebra with central charge $c=3k$.
The conformal boundary profile of the dilaton field does not modify the Virasoro algebra obtained under Drinfeld–Sokolov boundary conditions. In the conventional BF formulation, the ASA is generated by gauge transformations of the $\mathfrak{sl}(2,\mathbb{R})$-valued connection $a_{t}$, and this algebraic structure, including its central charge, remains intact. However, once the dilaton field $x$ is promoted to a dynamical boundary degree of freedom, the set of background-preserving residual gauge transformations becomes restricted. This restriction does not deform the Virasoro algebra itself, but it affects how the symmetry is realized on the boundary phase space.
Consequently, while the ASA remains a single copy of the Virasoro algebra with central charge $c=3k$, the boundary dynamics involves additional dilaton-dependent modes that are not fully captured by the Schwarzian action alone. In this sense, the dilaton plays a dual role: it determines the background geometry and constrains the admissible residual symmetries without altering the underlying algebraic structure.
@@
The analysis presented in this section is not in contradiction with the framework developed in Ref. Cardenas:2018krd , but rather aims to revisit closely related physical structures from a complementary conceptual and structural perspective. In particular, the notion of “asymptotic symmetry breaking” employed throughout this work should not be identified with the symmetry reduction arising from holonomy conditions, regularity requirements, or Casimir fixing in the sense discussed in Ref. Cardenas:2018krd . Instead, the emphasis here is placed on the restriction of admissible residual gauge transformations induced by the dynamical behavior of the dilaton supermultiplet.
In this sense, the term “breaking” does not refer to a violation of the underlying ASA, but to an algebraic and dynamical selection of stabilizers within the BF–theoretic formulation. As in Ref. Cardenas:2018krd , affine and superconformal boundary conditions are initially preserved. However, the focus of the present analysis lies on identifying a restrictive mechanism that already emerges at the level of bulk residual gauge symmetry, prior to imposing holonomy or additional regularity conditions. Accordingly, this section does not aim to reproduce the boundary dynamics derived in Ref. Cardenas:2018krd , but rather to elucidate the BF–theoretic origin of the associated asymptotic symmetry structure and to clarify the role played by the dilaton supermultiplet in shaping the physically admissible phase space.
@@
Section 4.1 defines a regime in which no highest-weight constraints are imposed on the state-dependent functions. The resulting asymptotic symmetry algebra is the full $\mathcal{N}=1$ affine $\mathfrak{osp}(1|2)_{k}$ algebra. In contrast, Section 4.2 introduces Drinfeld–Sokolov boundary conditions, which explicitly implement highest-weight constraints and thereby reduce the affine algebra to the classical $\mathcal{N}=1$ superconformal algebra. These two regimes correspond to distinct realizations of the boundary phase space and its associated symmetry structure.
@@ 4.1 $\mathcal{N}=1$ affine boundary conditions
The affine boundary conditions considered in this subsection are the supersymmetric counterpart of the bosonic affine setup discussed in Section 3.3 for $\mathfrak{sl}(2,\mathbb{R})$ JT dilaton gravity. In particular, the asymptotic symmetry is initially described by the affine $\mathfrak{osp}(1|2)_{k}$ algebra, while the dilaton supermultiplet dynamically restricts the set of admissible background-preserving residual gauge transformations.
The goal of this section is to construct an extended higher–spin $AdS_{2}$ framework by formulating it as an $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ dilaton supergravity theory defined with affine boundary conditions. Our analysis aims to determine the resulting ASA under the most general boundary conditions consistent with this setup.
Following the discussion in the previous section, we adopt the principal embedding of $\mathfrak{sl}(2,\mathbb{R})$ into $\mathfrak{osp}(1|2)$ as a subalgebra. This embedding provides a natural foundation for implementing affine boundary conditions in asymptotically $AdS_{2}$ spacetimes. To proceed with this construction, we introduce the gauge connection in the following form:
@@
In this framework, we obtain eight additional state-dependent functions, ${\cal X}^{i}$ and ${\cal Y}^{i}$, arising from the dilaton field expansion. Our goal is to derive the ASA associated with these affine boundary conditions via a canonical analysis. To this end, we systematically investigate all gauge transformations that preserve the structure of the boundary conditions defined above.
@@
As a final step in this construction, we define the canonical boundary charge $\mathcal{Q}_{a,x}[\lambda]$, which generates the gauge transformations presented in Eqs. (51)–(56). The variation of this charge, which encodes the structure of the ASA, is given by Banados:1994tn :
@@
Having established both the infinitesimal gauge transformations and the canonical boundary charge, we are now in a position to derive the ASA using the standard procedure outlined in Blagojevic:2002du . This is accomplished through the fundamental relation:
@@
Here, $\eta^{AB}$ denotes the invariant trace matrix, while $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ represent the structure constants of the $\mathfrak{osp}(1|2)$ algebra, with indices $(A,B=0,\pm 1,\pm\tfrac{1}{2})$. These constants satisfy the relation $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Under the most general (least restrictive) boundary conditions, the ASA of $\mathfrak{osp}(1|2)$ dilaton supergravity is characterized by a single copy of the affine $\mathfrak{osp}(1|2)_{k}$ algebra.
In the context of $\mathfrak{osp}(1|2)$ JT dilaton supergravity with affine boundary conditions, the ASA is given by a single copy of the infinite–dimensional affine $\mathfrak{osp}(1|2)_{k}$ algebra. The affine $\mathfrak{osp}(1|2)_{k}$ algebra remains unchanged.
The role of the dilaton supermultiplet is not to deform the affine algebra itself, but to influence the realization of its symmetries on the boundary phase space. Once a specific time–dependent dilaton background is specified, only those residual gauge transformations that preserve this background remain admissible. In other words, the set of realized asymptotic symmetries is dynamically reduced to the stabilizer subalgebra of the chosen dilaton configuration.
In this work, the term “asymptotic symmetry breaking” refers precisely to this reduction of admissible residual gauge transformations. It does not signify a deformation of the affine $\mathfrak{osp}(1|2)_{k}$ algebra, nor a modification of its structure constants or central extension. The realized symmetry is determined by the background-preserving residual gauge transformations selected by the dynamical dilaton supermultiplet.
This mechanism is structurally analogous to the symmetry reduction observed in the bosonic $\mathfrak{sl}(3,\mathbb{R})$ extension of JT gravity Ozer:2025bpb , where the affine algebra remains intact but the realized symmetry is reduced to the stabilizer subalgebra selected by the dilaton background.
The symmetry reduction originates from constraints imposed by the dilaton on the boundary phase space. Consequently, the residual symmetries encoded in the fields $a_{t}$ and $x$ correspond to boundary diffeomorphisms generated, respectively, by the bosonic and fermionic components $\epsilon^{i}$ and $\eta^{i}$ of the gauge parameter.
@@ 4.2 $\mathcal{N}=1$ superconformal boundary conditions
The superconformal boundary conditions imposed in this subsection provide the supersymmetric analogue of the bosonic superconformal (Drinfeld–Sokolov) reduction discussed in Section 3.4. Accordingly, the emergence of the classical $\mathcal{N}=1$ superconformal algebra originates from the imposed highest-weight (DS) constraints on the boundary data, rather than from a dilaton-induced dynamical selection.
In contrast to the affine regime discussed in Section 4.1, we now impose Drinfeld–Sokolov (highest weight) boundary conditions on the $\mathfrak{osp}(1|2)$-valued connection. This gauge choice implements an explicit algebraic reduction of the affine $\mathfrak{osp}(1|2)_{k}$ symmetry to the classical $\mathcal{N}=1$ superconformal algebra.
Unlike the dynamical symmetry restriction induced by the dilaton supermultiplet in the affine case, the present reduction originates directly from the boundary constraints themselves. Consequently, the emergence of the super-Virasoro structure is a structural consequence of the imposed highest-weight gauge.
In this section, we aim to analyze the ASA associated with the Brown–Henneaux boundary conditions in the context of $\mathfrak{osp}(1|2)$ supergravity. To this end, we impose the Drinfeld–Sokolov highest weight gauge condition on the $\mathfrak{osp}(1|2)$-valued connection defined in Eq. (48), which further restricts the coefficient structure. This gauge fixing yields the following field constraints:
@@
After implementing the above procedures, we are now positioned to derive the conformal ASA. As a consequence of the Drinfeld–Sokolov reduction, the residual gauge parameter $\lambda$ is effectively described by only four independent functions. These are denoted by $\epsilon\equiv\epsilon^{+1}$ and $\eta\equiv\eta^{+\frac{1}{2}}$, and are explicitly given by:
@@
Importantly, these gauge transformations provide key insights into the structure of the ASA Banados:1998pi . By analyzing the behavior of these asymptotic symmetries, one can determine the integral form of the variation of the canonical boundary charges,
@@
Tablo 1: superconformal spins for superconformal charges and related dilatons.
In the context of $\mathfrak{sl}(2,\mathbb{R})$ JT dilaton gravity, as well as in its $\mathfrak{osp}(1|2)$ extension under conformal boundary conditions, the ASA at the boundary is initially identified with the infinite–dimensional $\mathcal{N}=1$ superconformal algebra. However, the presence of a time–dependent dilaton field alters the algebraic structure of the asymptotic symmetry. As a result, even the expected preservation of the finite $\tt{O}\tt{S}p(1|2)$ subgroup within the full $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ symmetry is dynamically reduced, and the remaining infinite–dimensional symmetry components are broken.
@@
This choice retains only the $\mathcal{X}^{0}$ component while suppressing all fermionic modes $\mathcal{Y}^{\pm\frac{1}{2}}$ and the bosonic charges $\mathcal{X}^{\pm 1}$. Under the gauge transformations generated by the full affine algebra (Eqs. (51)–(56)), one finds that such a configuration leads to a restricted set of residual symmetries. In particular, the variations $\delta_{\lambda}\mathcal{Y}^{\pm\frac{1}{2}}$ and $\delta_{\lambda}\mathcal{X}^{\pm 1}$ vanish identically, while $\delta_{\lambda}\mathcal{X}^{0}$ survives only for $\epsilon^{0}$ and a subset of $\eta$–parameters. As a consequence, the boundary charges associated with broken modes cease to transform non–trivially and are effectively projected out of the ASA. This demonstrates, at a minimal level, how specific dilaton backgrounds can dynamically suppress certain symmetry generators, resulting in a reduced realization of the full $\mathfrak{osp}(1|2)_{k}$ structure. This kind of toy model construction, where a specific field configuration is used to analyze residual symmetry algebras, is also employed in the context of AdS2 boundary conditions Grumiller:2017qao .
@@
We therefore identify the ASA as the classical extended $\mathcal{N}=1$ superconformal algebra, rather than its quantum counterpart. This classical algebra arises naturally in our model and is consistent with the symmetry–breaking mechanism induced by the dilaton. Up to this point, our discussion has remained within the classical regime of the extended $\mathcal{N}=1$ superconformal symmetry. The quantum version of the theory, however, has already been presented in Eqs. (4)–(10). It is natural to ask whether the quantum theory can also be formulated in terms of a classical extended superconformal symmetry, or whether it necessitates a distinct quantum realization. While this remains a nontrivial question, preliminary evidence suggests that the quantum structure may be realized consistently, preserving the essential features of the classical algebra.